CAT Chat Minutes

October 4, 1996

 

Issues that require action and/or follow-up
1. How are the ASD studies progressing? Can the User Beam operating conditions be duplicated during the studies?

The studies have concentrated on the beam position monitors and the orbit correction systems. Progress is being made on both fronts. During the studies, the ID gaps are open, so BPLD trips only indicate but do not dump the stored beam. Except for runs at lower current, the storage ring studies conditions are duplicated.

2. What can you tell us about the downtime on Monday?

A linac klystron was failing and was not expected to last much longer. The decision was made to schedule its replacement after filling the storage ring on Monday morning. The linac downtime was expected to last 6-8 hours. However, a unexpected vacuum leak the in the linac waveguide delayed high voltage conditioning of the klystron and effectively ended the user run after the stored beam was lost at 1200 hours.

3. You want the ring commissioned, and we want stability. Have you thought about taking some additional studies time for dealing with stability improvements?

We need to continue installing FEs to meet the CAT schedules, and we need to improve the beam stability to meet the accelerator commissioning goals. The current apportionment of time reflects the best compromise for achieving both goals in a timely manner.

4. Both truck lock doors continue to be opened during deliveries. This causes temperature fluctuations in the hall. The problem will get worse in cold weather.

This will be brought to the attention of the appropriate people.

5. To diagnose a FE EPS trip, the EPS personnel had to drag a computer over to the equipment. This does not appear to be very efficient. Will this be the normal mode of operation?

A link from the EPS and PSS to EPICS is part of the current design. The implementation has started but will take several months to complete at all of the beamlines. This will allow floor coordinators and users to quickly identify the cause for a trip.

6. The WWW APS phone directory is still not available to the users.

Mohan Ramanathan will follow through on this item.

Information/Follow-up items
APS Operational Status Update:  Monday at 0800 hours User Beam will resume with a fill to 100 mA. No shielding verification is scheduled for next week. Thursday, October 10, an access into the storage ring will be required to investigate flange heating and vacuum anomalies in the front ends of 13ID and 3ID. If the beam goes down earlier in the week, we will attempt the access at that time and then would not need one on Thursday. If the current plan changes, we will notify you by CAT NET.